### v1

# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

Abstract
Distributed machine learning in high end-to-end latency and low, varying bandwidth network environments undergoes severe throughput degradation. Due to its low communication requirements, distributed SGD (D-SGD) remains the mainstream optimizer in such challenging networks, but it still suffers from significant throughput reduction. To mitigate these limitations, existing approaches typically employ gradient compression and delayed aggregation to alleviate low bandwidth and high latency, respectively. To address both challenges simultaneously, these strategies are often combined, introducing a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and model convergence rate. To achieve the balance under varying bandwidth conditions, an adaptive policy is required to dynamically adjust these parameters. Unfortunately, existing works rely on static heuristic strategies due to the lack of theoretical guidance, which prevents them from achieving this goal. This study fills in this theoretical gap by introducing a new theoretical tool, decomposing the joint optimization problem into a traditional convergence rate analysis with multiple analyzable noise terms. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression on training performance, filling a critical gap in understanding how compressed and delayed gradients affect training. Furthermore, by integrating the convergence rate with a network-aware time minimization condition, we propose DeCo-SGD, which dynamically adjusts the compression ratio and staleness based on the real-time network condition and training task. DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
1 Introduction
Recent breakthroughs in large language models (LLMs) (e.g., GPT-4 gpt-4 , DeepSeek-v3 deepseek ) have been accompanied by an explosive growth in model parameters, leading to unprecedented communication demands for distributed training scalalaws2 ; ScalaLaws1 . State-of-the-art (SOTA) distributed training strategies flashattention ; megatron-lm2 are predominantly optimized for high-speed interconnects like PCIe 5.0 and NVLink nvidia , where bandwidth exceeds $10$ Gbps and latency remains under $10$ ms. Although these assumptions are reasonable in well-resourced data centers, they often fail to hold in dynamic network environments with high latency and low bandwidth jaghouar2024opendiloco . Unfortunately, such unfavorable network conditions are common in distributed training over wide area networks (WAN), such as federated learning dagctmc and cross-geographic distributed learning hsieh2017gaia . WAN-based distributed training lowers the barrier for LLM pre-training, empowering individual researchers and small labs while reducing the reliance on costly, specialized networking SIGCOMM24StellaTrain .
In these challenging dynamic network settings, distributed SGD (D-SGD) and its variants, such as FedAVG fedavg and D-QSGD cui2021slashing , remain the dominant optimization methods due to their communication-friendly nature. Despite its low communication requirements, D-SGD still suffers from significant throughput degradation under high-latency and low-bandwidth conditions. As illustrated in Fig. 1, when the bandwidth drops below $2$ Gbps and the latency exceeds $200$ ms, the throughput of D-SGD in training GPT-2 with four nodes reduces to $50\%$ of maximum achievable throughput [^1] The maximum achievable throughput is determined by the compute capacity of the machines, without network constraints.. {wrapfigure}t0.4 [Uncaptioned image]
[…]
Gradient compression and delayed aggregation are common strategies to improve the throughput of D-SGD in a lossy manner, but joint usage introduces a complex three-way trade-off. Gradient compression [^2] In this work, the gradient compression algorithm carries vanilla error-feedback (EF) stich2019error by default. EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy. Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope. reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem dgc . Delayed aggregation DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem. To overcome the problem of low-bandwidth and high-latency at the same time in WANs, we need to use these two strategies jointly, introducing intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain, as it is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work. (2) Time-to-iteration: The effect of staleness and compression ratio on communication-computation overlap is not fully understood. Specifically, the existence of a threshold compression ratio, below which further compression no longer improves per-iteration time, remains unexplored. This threshold varies dynamically with the bandwidth and staleness.
To achieve optimal time-to-accuracy, we analyze the above two dimensions separately. First, in iteration-to-accuracy, we propose a new analytical tool called the Nested Virtual Sequence. This method builds on the original virtual sequence by constructing multiple virtual queues, breaking down the complex noise term into multiple analyzable noise terms. With this tool, we are the first to derive the convergence rate of distributed SGD with delayed aggregation and gradient compression (DD-EF-SGD). This convergence rate establishes the iteration-to-accuracy relationship and reveals that the value of staleness exponentially worsens the negative effect of the gradient compression on the model convergence. Second, we use mathematical induction to model the relationship between the end-to-end time and the number of iterations in DD-EF-SGD. In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm (shown in Fig. 3) that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
In a word, the contributions of our work are summarized as below:
  • We propose the Nested Virtual Sequence, a theoretical tool to analyze complex distributed SGD variants, by reducing it to a standard D-SGD process and several analytical noise terms. We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that the delayed aggregation exponentially worsens the negative effect of the compression on the training process.
  • We mathematically establish an approximate relationship between end-to-end time and accuracy. Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively calculates the optimal compression ratio and staleness according to the training environments.
  • We apply DeCo-SGD  in challenging networks environments (average bandwidth < $1$Gbps, latency $\geq$ $100$ms) across diverse model architectures including CNN, GPT and ViT. Experimental results validate the great three-way trade-offs of our design, achieving up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and the SOTA static strategy, respectively.
2 Preliminaries
We specifically concentrate on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
Figure 2: The running timelines for D-SGD and D-SGD variants. Both D-SGD and D-EF-SGD are serial processes, with D-EF-SGD achieving a reduced transmission time. Similarly, DD-SGD and DD-EF-SGD operate serially; while DD-SGD maintains the same communication time as D-SGD, DD-EF-SGD reduces transmission time.
2.1 The Optimization Problem of DML
In the data-parallel paradigm, the training dataset is partitioned across $n$ worker nodes, each holding a subset of the data. The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as a weighted sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
$\underset{\mathbf{x}}{\min}\left[f(\mathbf{x}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})\right].$
2.2 D-SGD and Its Variants
2.2.1 D-SGD
At $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$. $\gamma$ is the stepsize and these local gradients are aggregated to update the global model, following $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. As shown in Fig. 2, D-SGD is a serial computation process. In this process, computation and communication take place sequentially in each iteration, and communication includes end-to-end latency and transmission time.
2.2.2 D-SGD with Gradient Compression and EF (D-EF-SGD)
D-EF-SGD stich2020communication enhances D-SGD with the sparsification gradient compression [^3] In this work, the sparsification compressor in D-EF-SGD is Top-$k$ by default, a well-performed compressor in many hybrid gradient compressors stc ; CocktailSGD . by incorporating EF mechanisms to compensate for sparsification compression errors. Sparsification compression is $\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed gradient and the true gradient. We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$, $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker, and the update rules for D-EF-SGD are as follows:
$\begin{gathered}\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}),\quad\quad\quad\mathbf{e}_{t+1}^{i}=\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}-\Delta_{t}^{i},\quad\quad\quad\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\Delta_{t}^{i}.\end{gathered}$
As shown in Fig. 2, D-EF-SGD retains a serial process, but reduces transmission time compared to D-SGD.
2.2.3 D-SGD with Delay Aggregation (DD-SGD)
DD-SGD DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error . $\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}.$ The delay aggregation does not change the communication time for each iteration. As shown in Fig. 2, DD-SGD parallelizes computation and communication by performing the $t$-th computation at the $t$-th iteration and sending $t$-th gradient, but the aggregation applies the gradient computed at $(t-\tau)$-th iteration. When combined with gradient compression (DD-EF-SGD), the transmission time is further reduced, as shown in the last part of Fig. 2.
2.3 Virtual Sequence
Virtual Sequence stich2018sparsified ; mania2017perturbed is an analytical tool that simplifies the convergence analysis of D-SGD variants like D-EF-SGD. It defines an auxiliary sequence, $\tilde{\mathbf{x}}_{t}$, which evolves according to an idealized update rule and serves as a reference for the actual iterates, $\mathbf{x}_{t}$, accounting for noise introduced by compression and delay. We define the noise term as $B_{t}$, then we have $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$. We generalize updates as $\mathbf{v}_{t}$, (i.e., $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t}$) and $B_{t}$ can be computed in $B_{t+1}=B_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\mathbf{v}_{t}$. Then we have $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\textbf{g}_{t}^{i}$. This can help theoretically analyze D-SGD variants, but fails in complex variants like DD-EF-SGD, because the noise term decoupled from DD-EF-SGD cannot be analyzed. A new method is needed to analyze these complex D-SGD variants.
3 Theoretical Analysis of DD-EF-SGD
We derive the convergence rate of DD-EF-SGD. Notation list and detailed proof are in the Appendix.
3.1 Regular Assumptions
We assume functions are $L$-smooth with the gradient noise of D-SGD presumed to exhibit zero mean and a variance of $\sigma^{2}$. We gauge data heterogeneity using constants $\zeta_{i}^{2}>0$, which cap the variance among the $n$ workers. In Theorem 2, we posit that objective functions are $\mu$-strongly convex. Detailed assumptions are shown below.
Assumption 1 ($L$-smoothness). We assume $L$-smoothness of $f_{i},\ i\in[n]$, that is for all $\mathbf{x},\ \textbf{y}\in\mathbb{R}^{d}$:
$\lVert\nabla f_{i}(\textbf{y})-\nabla f_{i}(\mathbf{x})\rVert\leq L\lVert\textbf{y}-\mathbf{x}\rVert.$ (1)
Assumption 2 ($\mu$-strongly convexity). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
$f_{i}(\mathbf{x})-f_{i}(\textbf{y})\geq\langle\nabla f_{i}(\textbf{y}),\mathbf{x}-\textbf{y}\rangle+\frac{\mu}{2}\lVert\nabla f_{i}(\mathbf{x})-\nabla f_{i}(\textbf{y})\rVert^{2}.$ (2)
Assumption 3 (Bounded gradient noise). We assume that we have access to stochastic gradient oracles $\mathbf{g}^{i}_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for each $f_{i},\ i\in[n]$. We define the stochastic gradient noise of $i$-th worker as $\xi^{i}$ and for simplicity, we only consider the instructive case of uniformly bounded $\xi^{i}$ for all $\mathbf{x}\in\mathbb{R}^{d},\ i\in[n]$:
$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (3)
Assumption 4 (Measurement of data heterogeneity). We measure data dissimilarity by constants $\zeta_{i}^{2}\geq 0$ that bound the variance across the $n$ nodes. We have:
$\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+\lVert\nabla f(\mathbf{x})\rVert^{2},\quad\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2},\quad\forall\mathbf{x}\in\mathbb{R}^{d},i\in[n].$ (4)
These assumptions are standard, following previous works cui2022infocom ; stich2020communication . For Assumption 3, we can assume that the $\sigma^{2}$ increases with the gradient norm hardthreshold . In this way, by replacing $\sigma$ with a new $\sigma^{\prime}$, one can obtain a tighter bound. Similarly in Assumption 4, the general version is $\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+Z\lVert\nabla f(\mathbf{x})\rVert^{2}$, and we take $Z=1$ by default. One can take a more general bound by adding $Z$. However, extending the analysis to cover these generalized assumptions is beyond the scope of this work.
3.2 Nested Virtual Sequences (NVS)
To solve the problem that the traditional Virtual Sequence method cannot analyze complex D-SGD variants like DD-EF-SGD. We propose NVS, which is composed of two parts. In the first part, we have
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t},\quad\quad\quad B_{t+1}=B_{t}+\tilde{\mathbf{v}}_{t}-\mathbf{v}_{t}.$ (5)
Compared with Virtual Sequence, we change the second term in Eq. 5 from the update $\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$ into $\tilde{\mathbf{v}}_{t}$. We define the first virtual sequence $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$ and observe that $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\mathbf{v}_{t}-B_{t}-\tilde{\mathbf{v}}_{t}+\mathbf{v}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$. In the second part, we use the virtual sequence framework based on the new items $\tilde{\mathbf{x}}_{t}$ and $\tilde{\mathbf{v}}_{t}$. We define the second noise term as $\tilde{B}_{t}$ and have:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}.$
Similarly, it is easy to get $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. In this way, we decouple DD-EF-SGD into a standard D-SGD process and two analyzable noise terms:
$\mathbf{v}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\mathbf{g}_{t-\tau}^{i}+e_{t-\tau}^{i}),\quad B_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}e_{t-\tau}^{i},$
$\tilde{\mathbf{v}}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (6)
3.3 Convergence Rate of DD-EF-SGD
Theorem 1 (Non-convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{1}{\epsilon}+\frac{\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$ (7)
iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total number of iterations, chosen at random uniformly.
Remark 1.
($\phi$ determining the convergence in the non-degradation condition) In typical DML scenarios, $\frac{\tau}{n}\leq 1$. As the training scale increases, $\tau$ tends to decrease due to the larger model requiring larger computation time, and $n$ tends to increase, resulting in an even smaller $\frac{\tau}{n}$. While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in DML CocktailSGD ; grace , leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$. Then the convergence rate is $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{\sqrt{\frac{\phi}{\delta}}+\sqrt{\tau}}{\epsilon})$. Similar to prior works dagctmc , the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$. The second term $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$ has a greater impact on the overall convergence rate compared to the third term.
Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs pre-training scenarios, local datasets are typically allocated centrally, resulting in low data heterogeneity $\zeta$. Moreover, due to the small batch sizes commonly used in LLM training (limited by the constraint of GPU memory), $\sigma$ tends to be large. Under such circumstances, $\frac{\zeta^{2}}{\delta}\ll\sigma^{2}$ and $\phi$ captures the convergence rate. While in Federated Learning (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$. In this work, we focus on the first scenario where $\phi$ is the key factor that determines the convergence of DD-EF-SGD. For the analysis in Federated Learning or training small models, simply replacing $\phi$ with $\phi^{\prime}$ is enough and we do not discuss this later in the paper.
Remark 2.
(Degradation condition) DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$). When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into DD-SGD. Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\sqrt{\tau}}{\epsilon})$, same with the convergence rate of DD-SGD stich2022sharper ; stich2019error . When $\tau=0$, DD-EF-SGD degrades to D-EF-SGD, and $\phi=\frac{1-\delta}{\delta}$. Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{L}{\delta\epsilon})$, same with the convergence rate of D-EF-SGD stich2020communication ; dagc .
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$ (8)
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
Remark 3.
Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of DD-SGD (under $\delta=1$) and D-EF-SGD (under $\tau=0$). In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{\sqrt{\frac{L\phi}{\delta}}+\sqrt{L\tau}}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.
4 Mathematical Modeling Time to Accuracy in DD-EF-SGD
[…]
4.1 Time to Iteration in DD-EF-SGD
Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}$ satisfies the following approximation:
$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot S_{g}/a}{\tau+1},\frac{\delta\cdot S_{g}}{a},T_{\text{comp}}\right\}},$ (9)
where the approximation error is upper bounded by $O(\frac{1}{t})$.
Remark 4 (The optimal compression ratio $\delta^{*}$).
We can minimize $T_{avg}$ to $T_{comp}$ by enlarging $\tau$ and choosing a small $\delta$ Let $\delta^{*}(\tau)=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$, we can obtain that $\delta\leq\delta^{*}(\tau)$. Given $\tau\geq\frac{b}{T_{comp}}$, $\delta^{*}=\delta^{*}(\tau)$ This is due to the fact that the more aggressive $\delta$ cannot get the performance gain, which is covered by the delay while leading to accuracy degradation. In this way, $\delta^{*}$ is the maximal value in the given range.
4.2 DeCo: Joint Optimization for Delay Staleness and Gradient Compression in DML
Figure 3: The high level design of DeCo-SGD. Our design aims to minimize the end-to-end time while achieving the target accuracy. DeCo-SGD adaptively adjusts the compression ratio $\delta$ and delay staleness $\tau$ based on the dynamic network conditions monitored by the worker.
Input: $S_{g}$, $a$, $b$, $T_{comp}$
Output: Optimal $\tau$, $\delta$ with minimal $\phi$
$\phi_{min}=\infty$
for $\tau\in\{\lceil\frac{b+S_{g}/a}{T_{comp}}\rceil,\cdots,1+\lceil\frac{b}{T_{comp}}\rceil,\lceil\frac{b}{T_{comp}}\rceil\}$  do
      $\delta=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$;
     /* Calculate the value of $\phi$/
      $\phi=\frac{1-\delta}{\delta\cdot(1-\frac{\delta}{2})^{\tau}}$;
     if $\phi\leq\phi_{min}$ then
           $\phi_{min}=\phi$;
           $\tau^{*},\delta^{*}=\tau,\delta$;
           end if
           end for
          Return $\tau^{*}$, $\delta^{*}$;
Algorithm 1 DeCo
Input: $n$, $\gamma$, $S_{g}$, $T_{comp}$, initial parameters $\textbf{x}_{0}$, initial local error $\textbf{e}_{0}^{i}=\textbf{0}_{d}$, the frequency to update DeCo $E$
Output: $\textbf{x}_{T}$
for $i=1,\ldots,n$ do
       if $t\mod E==1$ then
             /* Update $\tau$ and $\delta$ with DeCo per $E$ iterations */
             Get $a$, $b$ from the network;
             $\tau$, $\delta$ = DeCo($S_{g}$, $a$, $b$, $T_{comp}$);
       end if
      $\hat{\Delta}_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i})$;
       $\textbf{e}_{t+1-\tau}^{i}=\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i}-\hat{\Delta}_{t}^{i}$;
       $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\hat{\Delta}_{t}^{i}$;
end for
Return $\textbf{x}_{T}$;
Algorithm 2 DeCo-SGD
Our design goal is to get the minimal end-to-end time to achieve a specific accuracy, which is equivalent to solving the minimum of the binary function $\phi$. Since $\tau$ is discrete, the minimum of $\phi$ cannot be directly solved by the partial derivatives. We use the critical condition $T_{avg}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima. The mathematical description of this goal is
$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\quad\textit{s.t.}\ T_{avg}=T_{comp},\quad\tau\in\mathbb{N},\quad\delta\in\left(0,1\right].$ (10)
Combined with Remark 4, we can rewrite Eq. 10 into:
$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}\quad\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{S_{g}}{a}}{T_{comp}}\rceil\right].$ (11)
We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo. 1).
The implementation of DeCo in D-SGD is called DeCo-SGD (pseudo-code shown in Algo. 2 and design shown in Fig. 3). DeCo-SGD sets a frequency $E$ to update DeCo. $E$ describes the sensitivity of DeCo. To enable frequent adjustments of the compression ratio and delay staleness in response to a dynamic network environment, a smaller $E$ value can be adopted, such as $E=1$. In the normal case, a middle $E$ like $100$ is enough. DeCo-SGD incurs a negligible extra cost compared to vanilla DD-EF-SGD and the extra time complexity is $\mathcal{O}(\frac{T}{E})$. The increase in $n$ does not result in additional cost.
5 Evaluation Experiments
5.1 Experimental Settings
We evaluate our approach on diverse CV and NLP tasks using CNN lecun1998gradient , ViT vit , and GPT gpt2 models, including FashionMNIST xiao2017fashion , CIFAR-$10$, ImageNet imagenet , and Wikitext wikitext . Our baselines include D-SGD, Accordion accordion , DGA DGA , and CocktailSGD CocktailSGD . CocktailSGD is a SOTA static strategy that combines gradient compression and delayed aggregation using fixed parameters. If not specified, all compressors are Top-$k$ gradient compression dagctmc , except for CocktailSGD.
5.2 Comparison of Training Speed
As illustrated in Fig. 4, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD across all tasks. All experiments were conducted with $4$ worker nodes. Take ViT@ImageNet as an example, when the accuracy reaches 70%, DeCo-SGD achieves a 4.99×4.$9$ speedup444speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD. over D-SGD and a 1.34×1.34$\times$1.34 × speedup over CocktailSGD. At 80% accuracy, the speedup increases to 5.01×5.01$\times$5.01 × and 1.37×1.37$\times$1.37 ×, respectively. Compared with other baseline methods, the benefits of DeCo-SGD are particularly obvious, mainly for two reasons: (1) their parameter settings are suboptimal; (2) they cannot adapt to the dynamic network environment in real time. In contrast, DeCo-SGD utilizes theoretical guidance and runtime feedback to jointly optimize compression and latency, thereby achieving a more effective communication-computing trade-off throughout the training process.
Figure 4: Training time comparison across different model-dataset pairs.
5.3 Scalability Verification of DeCo-SGD
DeCo-SGD is inherently scalable, as the computational complexity of DeCo-SGD in finding $\delta^{*}$ and $\tau^{*}$ is theoretically independent of the number of worker nodes. Fig. 5 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from $n=4$ to $n=32$. All experiments were conducted under a fixed latency of $200$ms, with bandwidth fluctuating around $100$Mbps.
The results demonstrate that DeCo-SGD consistently maintains performance improvements as the number of nodes increases. Specifically, for GPT@Wikitext, DeCo-SGD achieves up to 3.77×3.$7$ and 1.21×1.21$\times$1.21 × speed-up over D-SGD and CocktailSGD, respectively, even at $n=32$. Similarly, for ViT@ImageNet, DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over CocktailSGD. These results highlight the adaptability of DeCo-SGD to large-scale and high-precision distributed training tasks.
Figure 5: Training time comparison of different methods when the worker node scales. DeCo-SGD demonstrates superior scalability and efficiency compared to other methods across both settings.
5.4 DeCo-SGD under Different Bandwidths and Latency
Table 1 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes. The numbers in the table denote the training time (in seconds) required by each method to achieve a target perplexity ($22$ for the GPT model) or accuracy ($80$% for the ViT model). The results clearly demonstrate that under all network conditions, DeCo-SGD consistently outperforms other methods, achieving the lowest training times. This highlights its robustness against bandwidth and latency variations as well as its superior communication efficiency in challenging network environments. For instance, DeCo-SGD achieves up to 5.02×5.02$\times$5.02 × and 1.30×1.30$\times$1.30 × speed-ups over D-SGD and CocktailSGD, respectively.
Table 1: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. Detailed numbers are in the Appendix.
Model @DataSet a(Gbps), b(s) D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $6396.95\pm 32.36$ $1955.28\pm 13.59$ $5506.78\pm 26.77$ $1525.21\pm 12.91$ $\mathbf{1306.29}\pm\mathbf{10.29}$
$0.5,0.1$ $1571.37\pm 11.13$ $829.06\pm 6.65$ $1346.34\pm 10.75$ $712.93\pm 5.38$ $\mathbf{642.79}\pm\mathbf{5.74}$
$0.1,1.0$ $7232.88\pm 35.62$ $2510.26\pm 22.73$ $5764.60\pm 29.69$ $1683.11\pm 14.66$ $\mathbf{1440.59}\pm\mathbf{11.87}$
$0.5,1.0$ $2363.73\pm 13.87$ $1325.12\pm 10.90$ $1562.93\pm 11.89$ $1132.63\pm 9.91$ $\mathbf{997.25}\pm\mathbf{7.92}$
ViT @ImageNet $0.1,0.1$ $1373.84\pm 10.58$ $440.14\pm 3.28$ $1183.68\pm 7.85$ $364.77\pm 1.02$ $\mathbf{283.06}\pm\mathbf{1.80}$
$0.5,0.1$ $446.20\pm 3.39$ $274.60\pm 2.21$ $385.60\pm 1.57$ $219.39\pm 0.98$ $\mathbf{169.13}\pm\mathbf{0.47}$
$0.1,1.0$ $1530.57\pm 9.89$ $604.67\pm 2.79$ $1287.03\pm 10.70$ $505.91\pm 2.56$ $\mathbf{397.16}\pm\mathbf{1.39}$
$0.5,1.0$ $631.15\pm 3.33$ $373.79\pm 1.92$ $444.38\pm 2.18$ $278.57\pm 1.35$ $\mathbf{235.95}\pm\mathbf{0.88}$
6 Related Work
Gradient Compression Strategy. Gradient compression methods are generally categorized into quantization alistarh2017qsgd ; bernstein2018signsgd ; seide20141 , sparsity aji2017sparse ; lin2017deep ; stich2018sparsified , hybrid approaches dryden2016communication ; jiang2018sketchml , and low-rank strategies vogels2019powersgd . Traditional schemes often rely on fixed compression parameters, which limit their adaptability to varying network and training conditions. To address this, recent works have introduced dynamic and adaptive strategies. DC2 abdelmoniem2021dc2 adjusts communication based on real-time network dynamics, while DAGC lu2023dagc assigns compression aggressiveness based on each worker’s data volume. However, neither of them considers end-to-end latency. FedCG jiang2023heterogeneity selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes. L-GreCo markov2024greco formulates the per-layer compression allocation as a constrained optimization problem solved via dynamic programming to balance model size and error. However, neither of the above two methods takes into account the dynamic changes of the network environment.
Asynchronism. FedBuff nguyen2022federated uses a method of buffering asynchronous aggregation to improve communication efficiency and model performance. However, the buffering mechanism may increase memory overhead and delay the incorporation of fresh gradients. Mishchenko et al.mishchenko2022asynchronous rigorously proved theoretically and validated empirically that asynchronous SGD can achieve convergence comparable to mini-batch SGD, even in the presence of arbitrary gradient delays. CO2 sun2024co2 introduces local update and single-step asynchronous communication mechanisms to achieve complete overlap between communication and computing, but it assumes homogeneous device capabilities. To further address heterogeneity in node speed, Shadowheart SGD tyurinshadowheart dynamically adjusts the local update frequency and communication frequency of each node based on the different computing and communication speeds of each working node, thereby achieving the optimal time complexity in an asynchronous setting.
7 Conclusion
In this work, we address the challenge of efficient distributed training in high-latency, low-bandwidth and varying bandwidth networks, where D-SGD suffers significant throughput degradation. While gradient compression and delayed aggregation can mitigate bandwidth and latency issues, respectively, they introduce a complex trade-off among compression ratio, staleness, and model accuracy. To achieve the optimal balance during the whole training, we mathematically model it from two perspectives: time-to-iteration and iteration-to-accuracy. For the former, we derive the threshold compression ratio that maximizes computation-communication overlap under the given staleness and network conditions. For the latter, we introduce NVS to derive the convergence rate of DD-EF-SGD. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression, filling the critical theoretical gap in DD-EF-SGD. Building on these insights, we propose DeCo-SGD, an adaptive optimizer that dynamically adjusts compression and staleness based on network conditions. Our design achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over the SOTA method. The code will be released upon publication.
Appendix A Notation List
Table 2: Notation list.
Notation Description
$n$ number of workers
$\mathbf{T}$ total number of iterations
$f_{i}(\cdot)$ the local loss function of worker $i$
$f(\cdot)$ the global loss function, i.e., $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$
$\delta$ compression ratio ($0<\delta\leq 1$)
$\mathbf{C}_{\delta}(\cdot)$ the sparsification compressor with compression ratio $\delta$
$\mathbf{x}^{i}_{t}$ the local model parameter in the $t$-th iteration of the $i$-th worker
$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
$\mathbf{g}_{t}$ the average of $\mathbf{g}^{i}_{t}$, i.e., $\mathbf{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}^{i}_{t}$
$\gamma$ the stepsize
$\tau$ the delayed staleness
$\mathbf{e}_{t}^{i}$ the local error term of the $i$-th worker at the $t$-th iteration
$L$ L-smoothness
$\mu$ $\mu$-strongly convexity if $f_{i}$ is strongly convex
$\zeta_{i}$ the data heterogenority of worker $i$
$\zeta$ the global data heterogenority, i.e., $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2}$
$\xi^{i}$ the stochastic gradient noise of worker $i$
$\sigma^{2}$ the upper variance bound of the stochastic gradient noise
$a$ the bandwidth (bits/s)
$b$ the network end-to-end latency (s)
$S_{g}$ the size of the gradient (bits)
$T_{comp}$ computation time per iteration (s)
$T_{avg}$ average end-to-end training time per iteration (s)
$T_{sum}$ total end-to-end training time (s)
$TS_{k}$ End time of the k-th computation(s)
$TM_{k}$ End time of the k-th transmission(s)
$TC_{k}$ End time of the k-th communication(s)
Appendix B Detailed proof
B.1 Technical Results
We list lemmas derived from other works here to help us complete the whole proof. Detailed proof of these lemmas can be found from the reference and we do not write here.
Lemma 1.
If $c_{1},c_{2}\in\mathbb{R}^{d}$ then the Jensen’s inequality is: For all $\rho>0$, we have
$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert a\rVert^{2}+(1+\rho^{-1})\lVert b\rVert^{2}.$ (12)
This can be written as:
$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert a\rVert^{2}+\rho^{-1}\lVert b\rVert^{2}.$ (13)
Lemma 2 (The nature of Top-$k$).
By definition, the Top-$k$ compressor $\mathbf{C}_{\delta}$ is a mapping that has the property $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$:
$\mathbb{E}_{\mathbf{C}_{\delta}}\lVert\mathbf{C}_{\delta}(\mathbf{x})-\mathbf{x}\rVert^{2}\leq(1-\delta)\lVert\mathbf{x}\rVert^{2}.$ (14)
Lemma 3 (Lemma 27 of the work [19]).
Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
$r_{t+1}\leq r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $B>0$, $C,D\geq 0$ and step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$. Then there exists a constant stepsize $\gamma\leq\frac{1}{E}$ such that
$\frac{B}{T+1}\sum_{t=0}^{T}s_{t}\leq\frac{Er_{0}}{T+1}+2D^{1/3}\left(\frac{r_{0}}{T+1}\right)^{2/3}+2\left(\frac{Cr_{0}}{T+1}\right)^{1/2}.$ (15)
Remark 5.
To ensure that the right hand side in Eq. 15 is less than $\epsilon>0$,
$T=\mathcal{O}\left(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon}\right)\cdot r_{0}$
steps are sufficient.
Lemma 4 (Lemma 25 of the work [19]).
Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
$r_{t+1}\leq(1-\min\{\gamma A,F\})r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$, and for parameter $0<F\leq 1$. Then there exists a constant step-size $\gamma\leq\frac{1}{E}$ such that
$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+\min\left\{A,\frac{F}{\gamma}\right\}r_{T+1}$ $\leq r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}(T+1)\right]$
$\quad+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}}$
for $w_{t}:=(1-\min\{\gamma A,F\})^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and
$\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$
Remark 6.
Lemma  4 establishes a bound of the order
$\tilde{\mathcal{O}}\left(r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
that decreases with $T$. To ensure that this expression is less than $\epsilon$,
$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}\log\frac{1}{\epsilon}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}+\frac{E}{A}\right)$
steps are sufficient.
B.2 Proof of Theorem 1
Lemma 5.
Let $\left\{\mathbf{x}_{t},\tilde{\mathbf{x}}_{t},\hat{\mathbf{x}}_{t},B_{t},\tilde{B}_{t},\mathbf{v}_{t},\tilde{\mathbf{v}}_{t}\right\}_{t\geq 0}$ be defined as in DD-EF-SGD using NVS, with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. If $\gamma\leq\frac{1}{4L}$, then we have
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|B_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
Proof. We begin with the definition of $\hat{\mathbf{x}}_{t+1}$ and have
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)$ $=\mathbb{E}f\left(\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right)$ (16)
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
The first line is due to $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. The second line is due to Eq. 1 and the third line is due to Eq. 3. We estimate $A_{1}$ as
$A_{1}=\gamma\mathbb{E}\langle-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\rangle$ (17)
$=\gamma(\mathbb{E}\left\langle\nabla f\left(\tilde{\mathbf{x}}_{t}\right)-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right)-\nabla f\left(\tilde{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle-\mathbb{E}\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{g}_{t}\rangle)$
$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}.$
The last line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=2$, $c_{2}=4$. By bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
Lemma 6.
Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{4L\sqrt{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{16L^{2}}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\tau\gamma^{2}(1+T)\sigma^{2}}{n}.$
Proof. This lemma follows from the Lemma 20 in the previous work [14].
Lemma 7.
Define $\{B_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{4L\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\mathbb{E}\sum_{t=0}^{T}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Proof.
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$
$+\gamma^{2}(1-\delta)((1+\rho^{-1})\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2}).$
This inequality is due to Lemma 2, Eq. 4, 3,12.
By unrolling the recurrence and let $\rho=\frac{2-\delta}{\delta}$, we can derive the upper bound:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]\leq\frac{1}{4L^{2}}\sum_{i=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2\gamma^{2}(1-\delta)(1+T)}{\delta(1-\frac{\delta}{2})^{\tau}}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Theorem 1 (Non-convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}L}{n\epsilon^{2}}+\frac{L\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{L}{\epsilon}+\frac{L\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}L}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total iterations, chosen at random uniformly.
Proof. Let $G_{t}:=\sum_{i=0}^{t}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{i}\right)\right\|^{2}$. According to above lemmas (Lemma 5, 6, 7), we have
$\frac{\gamma}{2}G_{T}\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+2\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
Divide both sides of the inequality by $4\gamma(1+T)$, we have:
$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+4\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$. To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})r_{0}$ iterations are sufficient, which completes the proof.
B.3 Proof of Theorem 2
For further discussion, let’s define $F_{t}:=\mathbb{E}f(\mathbf{x}_{t})-f^{*}$ and $X_{t}:=\mathbb{E}\lVert\hat{\mathbf{x}}_{t}-\mathbf{x}^{*}\rVert^{2}$.
Lemma 8.
Let $\left\{\mathbf{x}_{t},\tilde{\mathbf{x}}_{t},\hat{\mathbf{x}}_{t},B_{t},\tilde{B}_{t},\mathbf{v}_{t},\tilde{\mathbf{v}}_{t}\right\}_{t\geq 0}$ be defined as in DD-EF-SGD using NVS, with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex. If $\gamma\leq\frac{1}{4L}$, then we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+\gamma^{2}\sigma^{2}+6\gamma L(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
Proof.
$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-x^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
We estimate the term $A_{2}$ as
$A_{2}$ $=-2\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\mathbf{x}^{*}\right\rangle-\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\tilde{\mathbf{x}}_{t}-\hat{\mathbf{x}}_{t}\right\rangle-\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\right\rangle)$
$\leq-\gamma\mu\mathbb{E}\lVert\mathbf{x}_{t}-\mathbf{x}^{*}\rVert^{2}-2\gamma F_{t}+\gamma F_{t}+4\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+4\gamma L\mathbb{E}\lVert B_{t}\rVert^{2}$
$\leq-\frac{\gamma\mu}{2}X_{t}-\gamma F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
The second line is due to the reason that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})$
$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}f_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
This completes the proof.
Lemma 9.
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{8\sqrt{L\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
Proof. This lemma follows from the Lemma 20 in the previous work [14].
Lemma 10.
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $B_{t}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{16\sqrt{L\phi}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
proof. Denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and we have:
$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
$+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}]$
$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we complete the proof.
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
Proof. By substituting Lemma 5, 6 into Lemma 4, we can get
$X_{t}\leq w_{T}X_{0}-\frac{\gamma}{2}\sum_{i=0}^{T}w_{t}F_{t}+\frac{W_{T}\gamma^{2}\sigma^{2}}{n}+6\gamma L(\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2})$
$\leq w_{T}X_{0}-\frac{1}{8}\sum_{i=0}^{T}w_{t}F_{t}+W_{T}\gamma^{2}\sigma^{2}+6LW_{T}\gamma^{3}(\frac{\tau\sigma^{2}}{n}+\frac{4\phi\zeta^{2}}{\delta}+2\phi\sigma^{2}).$
According to Lemma 4, we complete the proof.
B.4 Proof of Theorem 3
Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}:=\frac{TC_{t}}{t}$ satisfies the following approximation:
$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\}},$ (18)
where the approximation error is upper bounded by $O(\frac{1}{t})$.
proof. Based on the assumption, the iterative formula of $TS_{k}$ derived from the relationship between computation and communication can be written as :
$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2...\quad.$ (19)
with initial values $TS_{0}=TM_{0}=0$ and $TC_{k}=0$ for all $k\leq 0$. The analysis is divided into four cases, each corresponding to a specific regime determined by the relative magnitudes of $T_{\text{comp}}$, $\frac{\delta S_{g}}{a}$, and $\tau$.
Case 1: Computation-Dominated Regime
Consider the regime where $T_{\text{comp}}>\frac{\delta S_{g}}{a}$ and $\tau T_{\text{comp}}>\frac{\delta S_{g}}{a}+b$. We claim that:
$\forall k\in\mathbb{N}^{+},\quad TS_{k}=kT_{\text{comp}}.$
To verify this, we define the inductive hypothesis:
$P(k):\quad\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\forall i=1,\ldots,k.$
The base case $P(1)$ can be directly verified from the recurrence relations. Now assume $P(k)$ holds for some $k\in\mathbb{N}^{+}$. We analyze the $(k+1)$-th step:
  • If $k\leq\tau$, then $TC_{k-\tau}=0\leq TS_{k}$;
  • Otherwise, by the inductive assumption, we have:
    $TC_{k-\tau}=(k-\tau)T_{\text{comp}}+\frac{\delta S_{g}}{a}+b\leq kT_{\text{comp}}=TS_{k}.$
Thus, it follows that:
$TS_{k+1}=T_{\text{comp}}+\max\{TC_{k-\tau},TS_{k}\}=T_{\text{comp}}+TS_{k}=(k+1)T_{\text{comp}}.$
Moreover,
$TM_{k+1}=\frac{\delta S_{g}}{a}+\max\{TM_{k},TS_{k+1}\}=\frac{\delta S_{g}}{a}+TS_{k+1},$
since $TM_{k}\leq TS_{k+1}$ by the inductive assumption. Therefore, $P(k+1)$ holds, completing the induction.
It follows that:
$TC_{t}=TM_{t}+b=\frac{\delta S_{g}}{a}+tT_{\text{comp}}+b,$
and thus the approximation error satisfies:
$\left|TC_{t}-tT_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Case 2: Communication-Dominated Regime
Now consider the case where $\frac{\delta S_{g}}{a}>T_{\text{comp}}$ and $\tau\cdot\frac{\delta S_{g}}{a}>T_{\text{comp}}+b$. We claim that there exists $N_{1}\in\mathbb{N}^{+}$ such that:
$\forall k>N_{1},\quad TS_{k+1}-TS_{k}=\frac{\delta S_{g}}{a}.$
To establish this, we define the inductive statement:
$P(k):\quad\left\{\begin{aligned} TM_{i}&=T_{\text{comp}}+i\cdot\frac{\delta S_{g}}{a},\\ TM_{i}&\geq TS_{i+1},\end{aligned}\right.\quad\forall i=1,\ldots,k.$
For $i=1,\ldots,\tau+1$, direct computation gives:
$TS_{i}=iT_{\text{comp}},\quad TM_{i}=T_{\text{comp}}+i\cdot\frac{\delta S_{g}}{a},$
hence $P(\tau)$ holds. Assume $P(k)$ is true for some $k\geq\tau$.
Then:
$TM_{k+1}=\frac{\delta S_{g}}{a}+TM_{k}=T_{\text{comp}}+(k+1)\cdot\frac{\delta S_{g}}{a}.$
For the next computation step:
$TC_{k+1-\tau}+T_{\text{comp}}$ $=(k+1-\tau)\cdot\frac{\delta S_{g}}{a}+2T_{\text{comp}}+b$
$\leq(k+1)\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}\leq TM_{k+1},$
$TS_{k+1}+T_{\text{comp}}\leq TM_{k}+T_{\text{comp}}\leq TM_{k+1}.$
Therefore,
$TS_{k+2}=T_{\text{comp}}+\max\left\{TC_{k+1-\tau},TS_{k+1}\right\}\leq TM_{k+1},$
which implies $P(k+1)$ holds. By induction, $P(k)$ is valid for all $k\geq\tau$.
Consequently, for all sufficiently large $k$, the recurrence stabilizes and we obtain:
$TS_{k}=T_{\text{comp}}+TC_{k-\tau}=2T_{\text{comp}}+b+(k-\tau)\cdot\frac{\delta S_{g}}{a},$
with:
$\left|TC_{t}-t\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Finally, we verify the existence of $N_{1}>\tau$ such that $TC_{N_{1}-\tau}\geq TS_{N_{1}}$. If such $N_{1}$ does not exist, then both $\{TC_{k-\tau}\}$ and $\{TS_{k}\}$ would form arithmetic sequences with conflicting common differences, leading to a contradiction.
Case 3: Periodic Structure under Intermediate Delays
Consider the case where the computation time satisfies $T_{\text{comp}}>\frac{\delta S_{g}}{a}$ while the aggregated period satisfies $\tau T_{\text{comp}}\leq\frac{\delta S_{g}}{a}+b$. Under this condition, we can establish the existence of a positive integer $N$ such that the sequence $\left\{TS_{k+1}-TS_{k}\right\}_{k=N}^{T-1}$ becomes $(\tau+1)$-periodic. Specifically, for all $k\geq N$ and any $d\in\{1,\dots,\tau+1\}$, we have:
$TS_{(k+1)(\tau+1)+d}-TS_{k(\tau+1)+d}=T_{\text{comp}}+b+\frac{\delta S_{g}}{a}.$
This periodic property can be verified by explicitly analyzing the first $(2\tau+2)$ terms of the sequence $\{(TS_{k},TM_{k})\}$. We have:
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (20)
{$TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
Let us define the index group:
$A_{k}$ :=$:=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
$A_{k}+a$ :=$:=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
By rewriting the iterative update rule in the compact form $A_{k}=F(A_{k-1})$, it follows from translation invariance that:
$A_{k}+a=F(A_{k-1}+a).$
Now define the shifted sequence $A_{k}^{\prime}:=A_{k+1}-\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)$. Using Eq. (20), we obtain:
$A_{0}^{\prime}$ $=A_{1}-\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)=A_{0},$
$A_{1}^{\prime}$ $=F(A_{0}^{\prime})=A_{1}.$
Therefore, we conclude:
$A_{2}$ $=A_{1}^{\prime}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)=A_{1}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),$
and recursively,
$A_{k+1}=A_{k}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),\quad\forall k\in\mathbb{N}^{*}.$
Let $t=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
$\left|TC_{t}-t\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
$\leq\tau T_{\text{comp}}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
This confirms that $TC_{t}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
Case 4: Periodic Structure under Computation-Dominated Regime
In this case, we consider the regime where $T_{\text{comp}}<\frac{\delta S_{g}}{a}$ and $\tau\frac{\delta S_{g}}{a}\leq T_{\text{comp}}+b$. Under these conditions, we can establish that there exists a positive integer $N$ such that the update interval sequence $\left\{TS_{k+1}-TS_{k}\right\}_{k=N}^{T-1}$ exhibits $(\tau+1)$-periodicity. In particular, for all $k\geq N$ and for all $d\in\{1,\dots,\tau+1\}$, the following recurrence holds:
$TS_{(k+1)(\tau+1)+d}-TS_{k(\tau+1)+d}=T_{\text{comp}}+b+\frac{\delta S_{g}}{a}.$
Analogously to Case 3, we examine the first $(3\tau+3)$ elements of the sequence $\left\{(TS_{k},TM_{k})\right\}$ to validate the periodic structure:
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right.\quad$ $i=1,\dots,\tau+1;$
{$TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=\tau+2,\dots,2\tau+2;$
{$TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=2\tau+3,\dots,3\tau+3.$
We define the grouping $A_{k}$ as in Case 3. From the recurrence observed above, we find that:
$A_{2}=A_{1}+\left(T_{\text{comp}}+\frac{\delta S_{g}}{a}+b\right),$
and by translation invariance of the update dynamics, it follows that:
$A_{k+2}=A_{k+1}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),\quad\forall k\in\mathbb{N}^{*}.$
Thus, the $(\tau+1)$-periodicity is established, completing the proof for Case 4.
Now, let $t=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
$\left|TC_{t}-t\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
$\leq\tau\cdot\frac{\delta S_{g}}{a}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Summary of All Cases and Average Iteration Time
By integrating the results from all four cases, we obtain a unified expression for the asymptotic average iteration time:
$T^{\prime}_{\text{avg}}$ =$=\begin{cases}\dfrac{T_{\text{comp}}+\frac{\delta S_{g}}{a}+b}{\tau+1},&\tau<\tau_{\text{thres}},\\ \max\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},&\tau\geq\tau_{\text{thres}},\end{cases}\quad\text{with}\quad\left|TC_{t}-tT^{\prime}_{\text{avg}}\right|\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},$
where $T^{\prime}_{\text{avg}}:=\lim\limits_{k\to\infty}\frac{TS_{k}}{k}$ approximates the true average iteration time $T_{\text{avg}}$, and
$\tau_{\text{thres}}=\min\left\{\frac{\frac{\delta S_{g}}{a}+b}{T_{\text{comp}}},\frac{T_{\text{comp}}+b}{\frac{\delta S_{g}}{a}}\right\}.$
As a final conclusion, we estimate the average iteration time by:
$T_{\text{avg}}\approx T^{\prime}_{\text{avg}}=\max\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\},$
with the absolute error bounded by:
$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{t}=O\left(\frac{1}{t}\right).$
Appendix C Addendum to Evaluation Experiments
C.1 Experimental Environment
The experiments are conducted on an Ubuntu $24.04$ LTS server with an Intel Xeon Gold $6230$ CPU and $3$ Nvidia A40404040 GPUs, each with $48$GB memory. The Python version is $3.10.16$, and other used libraries are all based on the Python version. We use PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit. We use gloo as the communication backend.
C.2 Detailed Experimental Settings
Experiment tasks: The experimental setup encompasses three task types spanning computer vision (CV) and natural language processing (NLP), utilizing both traditional (CNN [29]) and state-of-the-art transformer-based models (ViT [30] and GPT [31]). The CNN architecture includes two convolutional layers followed by two fully connected layers, used widely in previous works [9, 20]. The ViT base model features approximately 86 million parameters, while the smallest GPT variant contains 124 million parameters.
To demonstrate the superiority of our design, we evaluate four representative tasks: FashionMNIST [32] and CIFAR-$10$ with CNN (denoted as CNN@FMNIST and CNN@CIFAR-$10$), ImageNet [33] using ViT (denoted as ViT@ImageNet) and Wikitext [34] on GPT (denoted as GPT@Wikitext). For CNN@FMNIST and CNN@CIFAR-$10$, the learning rate is $0.01$, and the batch size is $32$. For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$.
Baselines: We compare D-SGD, Accordion, DGA, and vanilla DD-EF-SGD as baselines. D-SGD is the case $\tau=0,\delta=1$, and While it does not suffer from accuracy degradation, it converges the slowest in high-latency and low-bandwidth scenarios when measured in terms of end-to-end time. Accordion [35] is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not [9]. The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients. The reason why we select Top-$k$ is its strong empirical performance, widespread adoption in DML, and straightforward implementation. DGA is a state-of-the-art algorithm originally designed for Federated Learning [18], which first introduces delay staleness to reduce the negative effect of the high latency on Federated Learning. We set $K=1$ in DGA to close the infrequent communication. CocktailSGD [21] uses the fixed $\tau$ and $\delta$. The hyperparameters are chosen by DeCo-SGD with $E=\infty$.
Network environment: We simulated challenging network environments (static high latency $\geq$ $100$ms, dynamic low bandwidth, average bandwidth $\leq$ $1$Gbps), where the dynamic volatility of the network can be referred to the following figure, which gives the network environment in the given environment (Section 5.2), and we discuss the superiority of the DeCo-SGD algorithm in different network conditions (Section 5.4).
C.3 Network Conditions of Section 5.2
As shown in Fig. LABEL:fig:network, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
C.4 Detailed Curves of Section 5.3 to Verify the Scalability of DeCo-SGD
The detail of Fig. 5 in the main text of the paper corresponds to Fig. 9 in the Appendix.
(a) ViT@ImageNet $n=4$
(b) ViT@ImageNet $n=8$
(c) ViT@ImageNet $n=16$
(d) ViT@ImageNet $n=32$
(e) GPT@Wikitext $n=4$
(f) GPT@Wikitext $n=8$
(g) GPT@Wikitext $n=16$
(h) GPT@Wikitext $n=32$
Figure 9: Convergence performance across different numbers of worker nodes ($n=4,8,16,32$) for five distributed optimization methods. Top row: Accuracy over time for ViT@ImageNet. Bottom row: Perplexity over time for GPT@Wikitext.
C.5 Detailed Results of Section 5.4
Fig. 10 in the Appendix provides a visual counterpart to the data presented in Table 1 in the main text. Table 3 extends more details in Table 1
Table 3: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. $\tau^{*}$ and $\delta^{*}$ are caculated by DeCo-SGD. The numbers in parentheses represent the speedup compared to DeCo-SGD.
Model @DataSet a(Gbps), b(s) $\tau^{*}$, $\delta^{*}$ D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $2,0.02$ $6396.95$ (4.90×4.90$\times$4.90 ×) $1955.28$ (1.50×1.50$\times$1.50 ×) $5506.78$ (4.22×4.$2$) $1525.21$ (1.17×1.17$\times$1.17 ×) $\mathbf{1306.29}$
$0.5,0.1$ $2,0.10$ $1571.37$ (2.44×2.$4$) $829.06$ (1.29×1.29$\times$1.29 ×) $1346.34$ (2.09×2.09$\times$2.09 ×) $712.93$ ($1\times1$) $\mathbf{642.79}$
$0.1,1.0$ $3,0.02$ $7232.88$ (5.02×5.02$\times$5.02 ×) $2510.26$ (1.74×1.74$\times$1.74 ×) $5764.60$ (4.00×4.$0$) $1683.11$ (1.17×1.17$\times$1.17 ×) $\mathbf{1440.59}$
$0.5,1.0$ $3,0.10$ $2363.73$ (2.37×2.37$\times$2.37 ×) $1325.12$ (1.33×1.$3$) $1562.93$ (1.57×1.57$\times$1.57 ×) $1132.63$ (1.14×1.14$\times$1.14 ×) $\mathbf{997.25}$
ViT @ImageNet $0.1,0.1$ $2,0.03$ $1373.84$ (4.85×4.85$\times$4.85 ×) $440.14$ (1.55×1.$5$) $1183.68$ (4.18×4.18$\times$4.18 ×) $364.77$ (1.29×1.29$\times$1.29 ×) $\mathbf{283.06}$
$0.5,0.1$ $2,0.14$ $446.20$ (2.64×2.64$\times$2.64 ×) $274.60$ (1.62×1.62$\times$1.62 ×) $385.60$ (2.28×2.28$\times$2.28 ×) $219.39$ (1.30×1.30$\times$1.30 ×) $\mathbf{169.13}$
$0.1,1.0$ $3,0.03$ $1530.57$ (3.85×3.85$\times$3.85 ×) $604.67$ (1.52×1.52$\times$1.52 ×) $1287.03$ (3.24×3.24$\times$3.24 ×) $505.91$ (1.27×1.27$\times$1.27 ×) $\mathbf{397.16}$
$0.5,1.0$ $3,0.14$ $631.15$ (2.67×2.67$\times$2.67 ×) $373.79$ (1.58×1.58$\times$1.58 ×) $444.38$ (1.88×1.$8$) $278.57$ (1.18×1.18$\times$1.18 ×) $\mathbf{235.95}$
(a) ViT@ImageNet bandwidth=$100$Mbps latency=$100$ms
(b) ViT@ImageNet bandwidth=$100$Mbps latency=$1000$ms
(c) ViT@ImageNet bandwidth=$500$Mbps latency=$100$ms
(d) ViT@ImageNet bandwidth=$500$Mbps latency=$1000$ms
(e) GPT@Wikitext bandwidth=$100$Mbps latency=$100$ms
(f) GPT@Wikitext bandwidth=$100$Mbps latency=$1000$ms
(g) GPT@Wikitext bandwidth=$500$Mbps latency=$100$ms
(h) GPT@Wikitext bandwidth=$500$Mbps latency=$1000$ms
Figure 10: Convergence performance across different numbers of worker nodes ($n=4,8,16,32$) for five distributed optimization methods. Top row: Accuracy over time for ViT@ImageNet. Bottom row: Perplexity over time for GPT@Wikitext.
Appendix D Others
D.1 UsedResources and Licenses
All external assets used in this paper are properly credited, and their licenses and terms of use have been explicitly acknowledged and respected. Specifically, we make use of several open-source resources as follows:
  1. 1.
    ViT model, released under the MIT License.
  2. 2.
    GPT2 model from Hugging Face, released under the MIT License.
  3. 3.
    docker-tc, used to simulate network conditions, released under the MIT License.
All of these assets were used strictly for research purposes in accordance with their respective licenses. We have cited the corresponding authors or repositories in our references where appropriate and ensured full compliance with their terms of use.
D.2 Limitations and Boarder Impact
Limitations: (1) Although DeCo-SGD currently supports only the D-SGD and its variants, we believe that D-SGD is applicable to a wide range of tasks, including NLP, CV and Objective Detection. If D-SGD is properly parameterized, its performance is no worse than popular optimizers in LLM training such as Adam [52]. (2) Device heterogeneity (different bandwidth/latency per node) is not considered, as current theoretical frameworks (based on convex optimization) cannot analyze such scenarios without prohibitive parameter search costs. Nevertheless, our approach already achieves $80\%$ throughput gains over SOTA optimizer [21] in dynamic networks, leaving limited room for further improvement via heterogeneity optimization.
Boarder Impact: Our work establishes theoretical foundations for LLM pre-training in low-bandwidth, high-latency networks. While employing lossy training techniques, we provide theoretical guarantees of training efficiency. This could significantly reduce the infrastructure costs for LLM pre-training, promoting broader participation in AI research and development.